Homework: HW-CH16 Score: 0 of 1 pt Problem 16.2 2 of 4 (1 complete) Given the following information about a product, at Michael Gibsons firm, what is the appropriate setup time? Setup labor cost$30.00 per hour Annual holding cost $29 per unit Daily production Annual demand Desired lot size 1,200 units/day 42,000 (280 days each xdaily demand of 150 units) 300 units Setup time-□ minutes (round your response to two decimal places).
Answer #1

Write the given information.

set up labor cost = $ 30.00 per hour

annual holding cost = $ 29 per unit

daily production = 1,200 units per day

annual demand = 42,000 (280 days each * daily demand of 150 units)

Daily Demand = 150 units

desired lot size = 300 units

S=[Q^2*H*(1-d/p)]/(2D)

Where

Q = desired lot size

H = Annual Holding Cost

d = Daily Demand

p= Daily production

D = Actual demand

S= Setup time

S = [300^2*29(1-150/1,200)]/2*42,000

= [600* 29 (1- 0.125)]/ 84,000

= [600 * 29 * 0.875]/84,000

= 15,225/84,000

= 0.18125 per hour

The setup labor cost would be divided by the set-up time

0.18125/ 30 = 0.0060416666 hours

0.0060416666 * 60 = 0.3625 mins